OpenStudy (anonymous):

Can I get a little help with this? f(-1)=4 f'(-1)=7 Let h(x)=(x^6)f(x). Evaluate h'(x) at x=-1

OpenStudy (anonymous):

By the Chain and Product Rule: h'(x) = (x^6)(f'(x)) + (6x^5)(f(x)) So, h'(-1) = ((-1)^6)(f'(-1)) + (-6)(f(-1) = (1)(7) + (-6)(4) = 7-24 = -17

OpenStudy (anonymous):

....ohhh. Thanks again!

OpenStudy (anonymous):

welcome.